OpenStudy (anonymous):

use the discriminant to determine the number of real - number solution for the equation:8x^2 + 8x +2 =0

OpenStudy (anonymous):

i do not know how to do this please help me.

OpenStudy (anonymous):

first, do you know what is the discriminant?

OpenStudy (campbell_st):

ok... so do you know the general quadratic formula..? \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

yap -b+- under square b^2 - 4ac/2a

OpenStudy (anonymous):

please

OpenStudy (anonymous):

hello

OpenStudy (campbell_st):

ok... so the conditions are discriminant > 0 you have 2 unequal zeros... if its a perfect square the roots are rational... e.g. x^2 + 5x + 6 = (x + 3)(x + 2) discriminant = 0 you have 2 equal roots ...e.g. x^2 + 6x + 9 = (x +3)(x +3) = (x +3)^2 and lastly discriminant < 0 you have complex roots or unreal roots... so in your question a = 8, b = 8 and c = 2 substitute them into the discriminant and see what you get... then compare it to the choices..

OpenStudy (campbell_st):

and use Discriminant = b^2 - 4ac

OpenStudy (campbell_st):

and one thing you need to know... every quadratic has 2 solutions... even though they may be equal.